A box contains 20 balls that are exactly alike except that 10 are red, 6 are blue, and 4 are black. (a) If a single ball is drawn from a box, what is the probability tat it is red? (b) If a single ball is drawn from the box, what is the probability that it is blue or black? (c) If instead two balls are drawn from the box at the same time, what would be the probability that both are black?


1
Expert's answer
2022-03-17T19:04:46-0400

a) "P={\frac {10} {20}}={\frac 1 2}"

b) "P={\frac {4+6} {20}}={\frac 12 }"

c) "P={\frac {4 \choose 2} {20 \choose 2}}={\frac 6 {190}}={\frac 3 {95}}"